| <p>23:00 DAQ restored and data now being written okay. Checks all okay, see screenshots below. Histograms reset. 40 bars in bottle @ 23:45 - should last 30 hours.</p>
<p>12:00 Checks all ok. No problems with aidas or writing to disk.</p>
<p>01:10 Checks all ok.</p>
<p>02:05 Checks all ok. See attatchment 6 for peak update after reset histograms 3 hours prior. Target okay, dropped a little to 6e12 but seems to be going up itself. Will monitor and then move position if needed.</p>
<p>03:10 During checks, noticed detector 4 had tripped so turned it off. Writing to disk increased alot, statistics increased to alost 250000. Put aida13-16 in low/medium energy. Started new run. (see attatchment 7-14 for before and afters)</p>
<p>RUN 91</p>
<p>03:15 Checks all ok after turning off detector 4. Moved target around and got abit more density but x is very sensitive and y did not seem to do much. Density 5.5e12.</p>
<p>04:30 Checks all ok. Target decreasing in density - monitoring it.</p>
<p>04:35 Target density back up to 1e13 after moving in y.</p>
<p>05:30 Checks all ok. Target density decreased slightly to 8e12</p>
<p>06:30 Checks all ok, target density stable for past hour</p>
<p>07:30 Checks all ok</p>
<p>08:30 Checks all ok, target density at 1e13 for past two hours</p>
<p>09:00 Run stopped</p>
<p> </p> |